Write the equation of a line parallel to 2x + 3y =7 and contains the point ( -9,4 )
Answer by Cromlix(4381)

First sort the equation into the
y = mx + c form
2x + 3y = 7
3y = -2x + 7
y = -2/3 x + 7/3
The gradient of this equation = -2/3
So a line parallel would have the same gradient.
Using y - b = m(x - a) and point (-9,4)
y - 4 = -2/3(x + 9)
y = -2/3x - 6 + 4
y = -2/3x - 2
or
3y = -2x - 6